OpenStudy (jas.2002):

Solve the equation for y. 5x – 7y = 10

pooja195 (pooja195):

Start by subtracting 5x from both sides \[\huge~\rm~5x -5x– 7y = 10-5x\]

pooja195 (pooja195):

What are you left with?

OpenStudy (jas.2002):

5x? I really don't know...

pooja195 (pooja195):

We subtracted 5x from both sides \[\huge~\rm~5x -5x– 7y = 10-5x\] \[\huge~\rm~– 7y = 10-5x\] Now just divide everything by -7

OpenStudy (jas.2002):

what do we divide by 7?

pooja195 (pooja195):

-7/-7=y 10/7=cannot be simplified -5/-7=5/7x a negative divided by a negative is a positive So y=____x+_____

OpenStudy (jas.2002):

so y= 5x + 7?Idk..

pooja195 (pooja195):

Remember we divided everything by -7

OpenStudy (♪chibiterasu):

When you don't end up with a whole number, you can just leave the right side as a binomial and put it over 7 to avoid the complication of decimals.

OpenStudy (jas.2002):

so is the answer -5x - 7?

OpenStudy (♪chibiterasu):

i.e. \[\frac{ x + y }{ z }\]

pooja195 (pooja195):

No We subtracted 5x from both sides \[\huge~\rm~5x -5x– 7y = 10-5x\] \[\huge~\rm~– 7y = 10-5x\] Now just divide everything by -7 \[\huge~\rm~-7y/-7=\frac{ -5 }{ -7 }x +\frac{ 10 }{ -7 }\]

pooja195 (pooja195):

Look at the last step. I divided everything by -7

OpenStudy (♪chibiterasu):

It would be -5x/-7 the variable isnt being multiplied into -7

OpenStudy (♪chibiterasu):

Ok yeah I guess it was jsut a typo then :x

OpenStudy (jas.2002):

ok..i think i got it.

pooja195 (pooja195):

What did you get?

pooja195 (pooja195):

Nope

OpenStudy (♪chibiterasu):

When dividing, the number you are dividing into goes on the bottom, not the top

pooja195 (pooja195):

A negative divided by a negative is a positive. So just 5/7x

OpenStudy (jas.2002):

ok thank you!

pooja195 (pooja195):

And you flipped the y intercept. -10/7

pooja195 (pooja195):

Put them together and make your line :)